Tohban Report 2015-08-19
From RHESSI Wiki
|Start Date:||12 Aug 2015|
|End Date:||19 Aug 2015|
|Next Tohban:||Lindsay Glesener|
|List all reports|
Solar activity was very low this week with a GOES base level to be at about B2; starting to raise on 2015-08-19. There are currently 4 ARs on the disk. AR12403 is on the east limb and shows some magnetic complexity; potentially producing some activity this week.
How many GOES flares occurred?
Flares above B, C, M, X class were 20 5 0 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 9 4 0 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 4 3 0 0
There were RHESSI flares/GOES flares 37 / 25 over the time range 12-Aug-15 19-Aug-15
SSR fill level between approx. 0 and 40%. We lost some downlink due to operational issues raising the minimum from 0% to about 20%. Despite low solar activity, we are only slowly gaining ground.
Cold plate temperatures rose from 129/130K to 132/133 as expected by orbital decrease of sunlight fraction.
GAPS IN APP_ID = 1 (VC1-SOH)
GAP START TIME GAP END TIME GAP (SEC) 2015-08-16T05:20:00.000 -- 2015-08-16T05:25:00.000 300.00000
No commanding has been performed this week. Detector performance appeared to be stable this week.
Three topics have been identified to possibly take action on:
Detector 9F livetime is currently at about 45%. The fast threshold could be in creased until a significant increase in livetime can be recorded.
Detector 4F and 6F shows elevated front slow and reset rates, but relatively low front fast rates.
Det. 4F slow threshold has been raised from 0x70 to 0x80.
2015-225-16:45:08 /IDPUTABLE4 OFFSET=FRONTSLOWDAC 2015-225-16:45:20 /IDPLOAD VALUE=0x80 ; was 0x70
The Att.3 threshold has not been adjusted yet due to the timing of BGS passes.
|Night time data (fronts)||+- 4 minutes|
|Night time data (rears)||+- 4 minutes|
|Require extra passes?||No|
|Requirement for moving pointer?||No|